OpenStudy (anonymous):

HELP! I need someone to explain how to solve this problem step-by-step. Solve the system by substitution -4.5x-2y=-12.5 3.25x-y=-0.75 Thanks!

OpenStudy (gorv):

okay so first we need to find value of x or y to substitute

OpenStudy (gorv):

pick any one equation

OpenStudy (gorv):

3.25-y=-0.75

OpenStudy (gorv):

add y on both side 3.25x=-0.75+y

OpenStudy (gorv):

add0.75 on both side

OpenStudy (gorv):

3.25x+0.75=y replace y by this value in other equation

OpenStudy (anonymous):

Ok, I think i get it from there. Would x=1?

OpenStudy (anonymous):

Would the final answer be (1,-17/2) @gorv ?